OpenStudy (anonymous):

No idea where to start here: "Express the following in terms of sin A, cos A or tan A where 0 degrees is less than or equal to A and A is less than or equal to 90 degrees - tan 120.5 degrees."

OpenStudy (anonymous):

Sorry I couldn't figure out to input an equation in the ask a question box

OpenStudy (anonymous):

Bit of a wordy question there indeed! So \(0 \le A \le 90\). Look at the unit circle, and you'll see that you have to look at the first quandrant of a circle.

OpenStudy (anonymous):

Yeah I know I have to use the circle, that's what we were learning about today. But what exactly do I do with it? What's the question asking?

OpenStudy (anonymous):

The number given is \(120.5^{o}\), which is in the second quadrant. That means your first step is to find the reference angle. Do you know how to do that bit?

OpenStudy (anonymous):

Hmm, is the reference angle the angle touching the middle of the circle?

OpenStudy (anonymous):

If you draw a line at \(120.5^{o}\), it makes another angle to the horizontal axis (x-axis). That is the angle you have to find.

OpenStudy (anonymous):

Isn't that just 180-120.5? 59.5?

OpenStudy (anonymous):

It is!

OpenStudy (anonymous):

Cool, but that's not the answer they're looking for is it? They want it expressed in terms or sin A cos A or tan A

OpenStudy (anonymous):

When you have "- \(tan(120.5)\) above, does the - mean negative, or is it a hyphen to separate the two parts of the question?

OpenStudy (anonymous):

We've found A so far, so we just have to decide whether to use sin, cos, or tan now.

OpenStudy (anonymous):

It's a separator

OpenStudy (anonymous):

Alright then. In the second quadrant, tan will be negative. By which I mean that \[tan(120.5) = -tan(59.5)\]

OpenStudy (anonymous):

We can also express \(tanA\) as \[\frac{sinA}{cosA}\]

OpenStudy (anonymous):

\[-tanA = \frac{-sinA}{-cosA}\]

OpenStudy (anonymous):

Sorry, ignore one of those negative signs

OpenStudy (anonymous):

Yeah, so how would you express the answer in the way the question says?

OpenStudy (anonymous):

You could do it in two ways.\[tan(120.5)=-tan(59.5)\] or\[tan(120.5)=-\frac{sin(59.5)}{cos(59.5)}\]

OpenStudy (anonymous):

Sweet, thanks. That's way easier than I thought. So just to make sure I've got it, I'll try and solve another example from the text book. \[\cos(135) = -\cos(25)\] Right? Or should I have not used cos?

OpenStudy (anonymous):

That's right, except that 180-135 = 45, not 25 :P

OpenStudy (anonymous):

Haha sorry about that. And \[-\sin (45) \] or \[tan(45)\] would still be right?

OpenStudy (anonymous):

Wait other way round. Switch the negative to the tan one

OpenStudy (anonymous):

You have to consider the relationship \[tanA=\frac{sinA}{cosA}\] If you want to get cos in terms of sin and tan, you have to rearrange the equation to get cos on its own. What would that look like?

OpenStudy (anonymous):

\[\sin(135) = cosA\] ?

OpenStudy (anonymous):

Never mind I think I've got it. Thanks for the help :)